# Are Colors Quanta of Light for Human Vision? A Quantum Cognition Study of Visual Perception

### v1

@@ 3 Quantum Measurement and the Bloch Model
To verify that quantum measurement gives rise to the mechanism of categorical perception, we must be able to express for states how much they differ from each other. But before we get to that, we must identify which states play the role of stimuli, and which states play the role of percepts in a quantum measurement process. The stimuli are represented by the states of the quantum entity independent of any measurement that would be going on, thus they are what in the quantum jargon are called the pure states. In Hilbert space, these pure states are represented by unit vectors of the Hilbert space. In the Bloch representation, it is the points of the Bloch sphere that represent the pure states of the considered quantum entity.
In what happens during a measurement as we represent it in the extended Bloch model, not only the places corresponding to points on the surface of the Bloch sphere play a role, but also places corresponding to points of the interior of the Bloch sphere are important, for example the point $A^{\prime}$ where the ball lands and sticks to the elastic after falling orthogonally on it. The points over which the elastic is stretched are also in the interior of the Bloch sphere. Besides the vector space formalism of quantum mechanics, where the main role is played by the complex Hilbert space, which is a vector space over the complex numbers, there also exists a density matrix formalism of quantum mechanics. However, both formalism have a fundamentally different mathematical structure, the Hilbert space is a vector space, as we already mentioned, while the set of density matrices is a convex space. Convex combinations of density matrices give another density matrix, while linear combinations of vectors give another vector. Also, both structures can live together on an underlying mathematical entity, e.g. density matrices are found in the Hilbert space formalism as density operators. Both are also found in the Bloch representation, the unit vectors of Hilbert space as the points of the surface of the Bloch sphere, and the density matrices as the points of the interior of the Bloch sphere. Let us introduce more systematically these mathematical structures by writing the different vectors with respect to the canonical basis.
$|1,0\rangle=(1,0)$ (21)
$|0,1\rangle=(0,1)$ (22)
We have
$|\theta,\phi\rangle=(\cos{\theta\over 2}e^{-i{\phi\over 2}},\sin{\theta\over 2}e^{i{\phi\over 2}})$ (23)
We already mentioned the density matrix formalism as also a quantum mechanical formalism. The points of the interior of the Bloch sphere correspond to such density matrices of this quantum formalism. We will denote a density operator corresponding to the point with spherical coordinates $(r,\theta,\phi)$, $\rho\in[0,1]$, $\theta\in[0,\pi]$, $\phi\in[0,2\pi]$ as $D_{(r,\theta,\phi)}$. Let us use some of the known properties to calculate the density matrix for some of the points of the Bloch sphere and of its interior. We already introduced the density matrix corresponding to a pure state in (12), hence in the notation we introduced this is $D_{(1,\theta,\phi)}$. Let us calculate the density matrices of the North and South Poles of the Bloch sphere, or of the points $A_{up}$ and $A_{down}$. We have
$D_{(1,0,\phi)}$ $=$ $\begin{pmatrix}\cos^{2}{\theta\over 2}&\cos{\theta\over 2}\sin{\theta\over 2}e^{-i\phi}\\ \cos{\theta\over 2}\sin{\theta\over 2}e^{i\phi}&\sin^{2}{\theta\over 2}\end{pmatrix}_{\theta=0,\phi}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$ (24)
$D_{(1,\pi,\phi)}$ $=$ $\begin{pmatrix}\cos^{2}{\theta\over 2}&\cos{\theta\over 2}\sin{\theta\over 2}e^{-i\phi}\\ \cos{\theta\over 2}\sin{\theta\over 2}e^{i\phi}&\sin^{2}{\theta\over 2}\end{pmatrix}_{\theta=0,\phi}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$ (25)
Let us now calculate the density operator $D_{A^{\prime}}$ representing the state of the entity when it is in the point $A^{\prime}$ as in Figure 2. We remark that $A^{\prime}$ lies on the line between $A_{down}$ and $A_{up}$ sticking on the elastic which is stretched between $A_{down}$ and $A_{up}$ on this line. Making use of a general property of the set of all density operators, i.e. that it is a set closed by convex combination, we know that $D_{A^{\prime}}$ is a convex combination of $D_{(1,\pi,\phi)}$ and $D_{(1,0,\phi)}$, which gives
$D_{A^{\prime}}=\lambda\begin{pmatrix}0&0\\ 0&1\end{pmatrix}+(1-\lambda)\begin{pmatrix}1&0\\ 0&0\end{pmatrix}=\begin{pmatrix}1-\lambda&0\\ 0&\lambda\end{pmatrix}$ (26)
for $\lambda\in[0,1]$. From Figure 2, given that $A^{\prime}$ is obtained by projecting orthogonally to the line between $A_{down}$ and $A_{up}$, we have that $A^{\prime}$ lies on the line between $A$ and the point with spherical coordinates $(1,\theta,\phi+\pi)$, to which corresponds the density operator
$D_{(1,\theta,\phi+\pi)}$ $=$ $\begin{pmatrix}\cos^{2}{\theta\over 2}&\cos{\theta\over 2}\sin{\theta\over 2}e^{-i(\phi+\pi)}\\ \cos{\theta\over 2}\sin{\theta\over 2}e^{i(\phi+\pi)}&\sin^{2}{\theta\over 2}\end{pmatrix}$ (27)
$=$ $\begin{pmatrix}\cos^{2}{\theta\over 2}&-\cos{\theta\over 2}\sin{\theta\over 2}e^{-i\phi}\\ -\cos{\theta\over 2}\sin{\theta\over 2}e^{i\phi}&\sin^{2}{\theta\over 2}\end{pmatrix}$
This means that we have
$D_{A^{\prime}}=\mu D_{(1,\theta,\phi)}+(1-\mu)D_{(1,\theta,\phi+\pi)}$ (28)
for $\mu\in[0,1]$. From (26) and (28) follows that we must have
$\mu\cos{\theta\over 2}\sin{\theta\over 2}e^{-i\phi}-(1-\mu)\cos{\theta\over 2}\sin{\theta\over 2}e^{-i\phi}=0$
$\Leftrightarrow\mu-(1-\mu)=0$
$\Leftrightarrow\mu={1\over 2}$ (29)
and
$\lambda=\sin^{2}{\theta\over 2}$ (30)
This gives us
$D_{A^{\prime}}$ $=$ $\begin{pmatrix}\cos^{2}{\theta\over 2}&0\\ 0&\sin^{2}{\theta\over 2}\end{pmatrix}=\cos^{2}{\theta\over 2}\begin{pmatrix}1&0\\ 0&0\end{pmatrix}+\sin^{2}{\theta\over 2}\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$ (31)
$=$ $\cos^{2}{\theta\over 2}D_{A_{up}}+\sin^{2}{\theta\over 2}D_{A_{down}}$
Let us note that it follows from (28) that we also know the density matrix of each point lying on the line between the two points of the surface of the Bloch sphere, the point $A$ with coordinates $(1,\theta,\phi)$ and the point with coordinates $(1,\theta,\phi+\pi)$, which is mirrored with respect to the North-South axis of the Bloch sphere. For a value of $\mu$ ranging from -1 to +1, the density matrix is given by
$D_{A^{\prime},\mu}$ $=$ $\mu D_{(1,\theta,\phi)}+(1-\mu)D_{(1,\theta,\phi+\pi)}$ (32)
$=$ $\begin{pmatrix}\cos^{2}{\theta\over 2}&\mu\cos{\theta\over 2}\sin{\theta\over 2}e^{-i(\phi+\pi)}\\ \mu\cos{\theta\over 2}\sin{\theta\over 2}e^{i(\phi+\pi)}&\sin^{2}{\theta\over 2}\end{pmatrix}$
the density matrix of the point corresponding to the value of $\mu$ on the line between the points $(1,\theta,\phi+\pi)$ and $(1,\theta,\phi)$. The parameter $\mu$ occurs only in the diagonal terms of the density matrix, and the orthogonal falling of the ball on the elastic takes place on the line connecting the two points $(1,\theta,\phi+\pi)$ and $(1,\theta,\phi)$, and that orthogonal falling occurs when $\mu$ goes from 1 to 0. In the jargon of quantum mechanics, the decreasing of the non-diagonal terms of the density matrix is named ‘decoherence,’ and is described as a phenomenon of decreasing quantum nature and increasing classical nature.
Let us briefly summarize the elements of a quantum measurement and its Bloch sphere representation. Each point $(r,\theta,\phi)$, $r\in[0,1]$, $\theta\in[0,\pi]$, $\phi\in[0,2\pi]$, within the sphere shown in Figure 2, corresponds to a density operator $D_{(r,\theta,\phi)}$, and the points of the sphere surface, i.e. $r=1$, correspond to pure states of the quantum entity. The points of the interior of the sphere correspond to density states of the quantum entity. With a measurement corresponds a convex subspace of the interior of the sphere, in the case of the measurement with possible final states after the measurement, $A_{up}$ and $A_{down}$, this convex subspace is given by the line segment from $(1,\pi,\phi)$ to $(1,0,\phi)$, where the elastic is stretched. But all lines passing through the center of the sphere in Figure 2 harbor such an equivalent measurement. The points on the line all correspond to a density state, except for the end points, which are indeed part of the surface of the sphere, and thus correspond to a pure state. When a measurement commences, on an entity that is in a pure state, such as this one in the point $A$, an expectation is allowed to play a role of what this state in the point $A$ means to the measuring apparatus located on a line between two diametrical points of the sphere surface, such as the points $A_{down}$ and $A_{up}$. That is the meaning of the orthogonal projection which brings the state in $A$, to a state in $A^{\prime}$. Once the state has become the one corresponding to the point $A^{\prime}$, then the entity is in a density state. The transformation from the state corresponding to $A$ to the state corresponding to $A^{\prime}$ can be read on the matrix representation, namely it is the non-diagonal terms of the matrix of the density state that disappear to arrive at the state corresponding to $A^{\prime}$.
4 Categorical Perception and Quantum Measurement
As we already mentioned, in this article we want to show that a quantum measurement structurally incorporates the mechanism of categorical perception. At the same time, we want to propose an interpretation of ‘what exactly could be imagined to happen during a quantum measurement’ with the intention of representing the ‘cognition aspect’ of this event as clearly as possible. In this sense, our work also provides additional evidence for the ‘conceptuality interpretation’ of quantum mechanics an original interpretation (Aerts, 2009b) that is being further developed in our Brussels research group (Aerts et al., 2020). Indeed, the basic hypothesis of the conceptuality interpretation is that a quantum measurement has the characteristics of a cognitive process, as a consequence the quantum entities which are the subject of the measurement have a conceptual nature. When we mention that the basic hypothesis of conceptuality interpretation consists in postulating the presence of a cognitive interaction in the micro world we do not mean that this is a cognitive interaction very similar to the human cognitive interaction, perhaps and even probably so, it is of an alien nature. We do mean that there are some basic aspects present that we also find in the human cognitive interaction, for example ‘the use of concepts as communication tools’ (Aerts, 2009b; Aerts et al., 2020). If the phenomenon of categorical perception can be shown to be present in what is occurring during a quantum measurement then this brings extra evidence for the hypothesis that a quantum measurement is a cognitive process. Ultimately, we also want to show that it makes sense to consider colors as the quanta of light for human visual perception similar to how photons are the quanta of light for a physical measurement device during a quantum measurement.
To inspire our interpretation of what could be taking place during a quantum measurement, we present an experiment conducted by Jerome Bruner and Postman, in which elements of human perception are brought out (Bruner & Postman, 1949). Note that Bruner, considered to be one of the important psychologists of the twentieth century, and Postman worked on perception in the 1940’s, an era where cognitive science was not yet defined as a sub discipline in psychology. In the experiment we will describe in some detail, Bruner and Postman’s intention was to confront participants with stimuli they called ‘incongruent’. To make clear what this term ‘incongruent’ means, we outline the view on perception in which Bruner and Postman framed the experiment. They worked within a view of ‘perception’ in which the hypothesis is advanced that a perception corresponding to a stimulus is an event in which there is, what they call, a ‘construction-defense’ balance at work. From a pattern of expectations, a structure present in the mind of the one who perceives, ‘construction’ takes place when a stimulus meets an expectation, while ‘defense’ takes place when a stimulus counters an expectation. In the experiment we will now describe, participants are confronted with stimuli that confirm but also contradict the structure of expectations with which participants’ minds take part in the experiment. The purpose of the experiment is to distinguish and analyze different phases of dealing with this contradiction, which Bruner and Postman name ‘incongruity’. We are particularly interested in one of the phases, which they call the ‘dominance reaction’. Let us now first describe the experiment in Bruner & Postman (1949) and then further on specify which aspects we will use to craft our interpretation of a quantum measurement.
Twenty-eight participants, students at Harvard and Radcliffe, were shown successively by tachistoscopic exposure five different playing cards. From one to four of these cards were incongruous – color and suit were reversed. Order of presentation of normal and incongruous cards was randomized. The normal and ‘trick’ cards used were the following. Normal cards (printed in their proper color) five of hearts, ace of hearts, five of spades, seven of spades. Trick cards (printed with color reversed), three of hearts (black), four of hearts (black), two of spades (red), six of spades (red), ace of diamonds (black), six of clubs (red). Fourteen orders of presentation were worked out, and two subjects were presented the cards in each of these orders. There were three types of stimulus series (1) a single trick card embedded in a series of four normal cards, (2) a single normal card embedded in a series of four trick cards, (3) mixed series in which trick and normal cards were in the ratio of 3 to 2 or 2 to 3. Each card was presented successively until correct recognition occurred, three times each at 10 ms, 30 ms, 50 ms, 70 ms, 100 ms, 150 ms, 200 ms, 250 ms, 300 ms, 400 ms, 450 ms, 500 ms, and then in steps of 100 ms to 1000 ms. If at 1000 ms recognition did not occur, the next card was presented. In determining thresholds, correct recognition was defined as two successive correct responses. At each exposure, the subject was asked to report everything he or she saw or thought he or she saw. The cards were mounted on a medium gray cardboard and were shown in a Dodge-Gerbrands tachistoscope. The pre-exposure field was of the same gray color and consistency as the exposure field save that it contained no playing card. The light in the tachistoscope was provided by two G E. daylight fluorescent tubes.
For those readers who would like more details about the experiment and its analysis and conclusions, we refer to (Bruner & Postman, 1949), we are particularly interested in the different reactions of the participants, and especially this one that Bruner and Postman called the ‘dominant reaction’. This dominant reaction to the cards shown in swift consists of providing a response that is entirely within the expectation pattern of the participant, namely that these are standard cards that have not been tricked. With respect to the tricked cards, this reaction consists, essentially, of a ‘perceptual denial’ of the incongruous elements in the stimulus pattern. Faced with a red six of spades, for example, a subject may report with considerable assurance, ‘the six of spades’ or the ‘six of hearts’, depending upon whether he is color or form bound. In the one case the form dominates and the color is assimilated to it, in the other the stimulus color dominates and form is assimilated to it. In both instances the perceptual resultant conforms with past expectations about the ‘normal’ nature of playing cards. We will not describe the other possible reactions further, as we are concerned with this one, incidentally, most common reaction. With longer times of showing the cards, and repeating the experiment with the same participants, a different pattern does develop, with the participants eventually finding out that tricked cards are being used – a real repetition of the experiment, to calculate e.g. the collapse probabilities, should, for this reason, always invite other participants.
In Bruner and Postman’s work, perception is considered a phenomenon that, in addition to the influence of the stimulus, is also determined by the presence of the expectation pattern of the one who perceives. True, the model of human perception proposed by them contains elements that are probably only relevant to the human specificity of what perception and cognition are. And, if as a basic hypothesis of the conceptuality interpretation of quantum mechanics the quantum measurement is considered as a cognitive process, in which then also a perceptual process takes place, then, like we mentioned already, it is not meant that the specific properties of human cognition and perception will also be present. In this sense, we are primarily interested in the fundamental and general properties of cognitive and perceptual processes, the formation of concepts itself probably being the most important one. The evidence gathered so far in support of the conceptuality interpretation is largely based on the hypothesis that quantum entities are conceptual entities, hence that the cognitive process makes use of ‘concepts’. The structure underpinning the phenomenon of categorical perception is already more specific, but still fundamental and general enough so that, given its abundance in human cognition, it makes sense to investigate its presence in a quantum measurement. Its identification in a quantum measurement will also allow us, as we will show, to distinguish precisely its most fundamental and general properties. But first, we want to point out some of these properties inspired by Bruner and Postman’s view of human perception.
According to Bruner and Postman, the set of percepts is primarily a set of ‘expectations related to the stimuli’. We will, for a quantum measurement, identify this set of percepts with the set of points on the centerline of the Bloch sphere, that is, where in our extended Bloch model the elastic is stretched. Where are the stimuli for a quantum measurement? The stimuli are assumed to be independent of any measurement and so, for a quantum measurement, they are the pure states of the entity in question, and within the Bloch sphere representation they are located in the points of the surface of the sphere. The orthogonal falling of the ball of a point of the sphere surface on the elastic band represents the interaction of the cognitive apparatus with the entity under consideration. This means that the transformation from a point on the surface of the Bloch sphere, say the point $A$, to a point on the centerline where the elastic is stretched, say the point $A^{\prime}$, represents the interaction between expectation, starting from the cognitive apparatus, on the one hand, and perception, starting from the considered entity, on the other.
If we want to give an expression to the mechanism of categorical perception, we must be able to estimate and/or measure distances between both pure states, the stimuli, located in points on the surface of the Bloch sphere, and density states, the percepts, located in the points of a centerline of the Bloch sphere. Using these distances, the differences between stimuli and the differences between percepts can be expressed quantitatively. However, we must be careful in choosing how we will measure distances in the Bloch sphere representation, since there is no unique obvious metric on the whole set of quantum states. One of the well-studied partial metrics on the set of quantum states is the trace distance
$T(D_{1},D_{2})$ $=$ ${1\over 2}Tr\left[\sqrt{(D_{1}-D_{2})^{*}(D_{1}-D_{2})}\right]$ (33)
@@ 4 Categorical Perception and Quantum Measurement
It can be shown that for a two-dimensional quantum entity, i.e., a qubit, modeled with the Bloch sphere, the trace distance is equal to half the Euclidean distance in the three-dimensional Euclidean space of which the Bloch sphere is a part. For two pure states $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, hence represented by two points of the surface of the Bloch sphere, it can be shown that the trace distance is given by
@@
That means, for example, that the trace distances between the quantum state where $A$ is and the North and South Pole of the Bloch sphere are given by
$T(D_{(1,\theta,\phi)},D_{(1,0,0)})$ $=$ $\sqrt{1-|\langle\theta,\phi|0,0\rangle|^{2}}=\sqrt{1-|\langle(\cos{\theta\over 2}e^{-i{\phi\over 2}},\sin{\theta\over 2}e^{i{\phi\over 2}})|(1,0)\rangle|^{2}}$ (35)
$=$ $\sqrt{1-\cos^{2}{\theta\over 2}}=\sin{\theta\over 2}$
$T(D_{(1,\theta,\phi)},D_{(1,\pi,0)})$ $=$ $\sqrt{1-|\langle\theta,\phi|\pi,0\rangle|^{2}}=\sqrt{1-|\langle(\cos{\theta\over 2}e^{-i{\phi\over 2}},\sin{\theta\over 2}e^{i{\phi\over 2}})|(0,1)\rangle|^{2}}$ (36)
$=$ $\sqrt{1-\sin^{2}{\theta\over 2}}=\cos{\theta\over 2}$
We can now see interesting connections using the simple geometry of the Bloch representation, let us consider Figure 3. It represents the Bloch representation with $\phi={\pi\over 2}$, hence $A$ laying in the $zy$-plane. The triangle $A_{up}$, $A$, $A_{down}$ is a right-angled triangle with right angle in $A$. The hypothenuse of this right-angled triangle is the centerline connecting $A_{up}$ to $A_{down}$ and it has length equal to $2$.
@@
The angle in $A_{down}$ of the triangle is equal to half the angle $\theta$, and so the length of the line segment connecting $A_{up}$ to $A$ is equal to $2\sin{\theta\over 2}$. In a similar manner, the figure shows us that the length of the line segment connecting $A$ and $A_{down}$ is given by $2\cos{\theta\over 2}$. This agrees with the calculation of the trace distance we made for the states connected to corresponding points of the Bloch sphere, as given in (35) and (36), reminding us that for a qubit the trace distance is equal to half the Euclidean distance. This reasoning and additional calculations, also the general expression for the trace distance between pure states, makes it clear that the trace distance is not the metric to be used between pure states. Indeed, any effect due to quantum coherence disappears when using this metric, since in the general expression (34) the only reference to the states is the inner product between the two states, and hence traces of quantum coherence effects are no longer present.
@@
There is also a purely geometric way by which we can see that the trace distance is unsuitable for measuring distances between pure states. Consider any two pure states, which gives us two points of the surface of the Bloch sphere. If we connect these two points with a line, we see that the points of the line different from these two points always lie in the interior of the Bloch sphere and thus correspond to density states. It is clear that defining a distance associated with a specific geometric entity, the straight line in the three-dimensional Euclidean space to which the Bloch sphere belongs, carries limitations when, as is the case with pure states, this distance is intrinsically intended for elements of a subset with a structure in which the straight line is not contained, and we mean here the spherical surface of the Bloch sphere, to which the pure states are restricted.
Figure 3: Using the simple geometry of the Bloch sphere, we can recover the trace distance we calculated in (35) and (36) as illustrated in this figure. The triangle $A_{up}$, $A$, $A_{down}$ is rectangular which makes its angle in $A_{down}$ equal to ${\theta\over 2}$, and consequently, knowing that the distance between $A_{up}$ and $A_{down}$ is equal to $2$, the figure shows us that the distance between $A_{up}$ and $A$ is equal to $2\sin{\theta\over 2}$, and the distance between $A_{down}$ and $A$ is equal to $2\cos{\theta\over 2}$. Since the trace distance is equal to half the Euclidean distance for the Bloch sphere of a qubit, we have derived the results of (35) and (36) from the geometry of the Bloch sphere.
@@
So we must ask ourselves at this point of our analysis which the proper metric to be used on pure quantum states is. A natural notion of distance between pure states should only account for the pure states, that are in-between the points on the Bloch sphere representing these states, the length of the circular arc connecting these points would be such a quantity. More precisely, if $\theta$ and $\alpha$ are the polar angles of the pure states $|\theta,\phi\rangle$ and $|\alpha,\phi\rangle$ (see Figure 4), the distance to be used to measure how much they are separated from each other, normalized to 1, would then be
$d_{pure}(|\theta,\phi\rangle,|\alpha,\phi\rangle)={1\over\pi}|\theta-\alpha|$ (37)
@@
Note that it is sufficient to define the distance for two pure states with equal azimuthal angle $\phi$. Indeed, we can always choose a North Pole and South Pole for any two pure states such that both states lie in the plane with equal azimuthal angle, so that then the arc of the circle through both points is determined by the difference of the polar angles of both states.
@@
This distance also results from the angle between pure states calculated from their Hilbert space inner product, which shows that it is a natural distance associated with the Bloch representation of pure states. The notion of fidelity
$F(D_{1},D_{2})=\left(tr\sqrt{\sqrt{D_{1}}D_{2}\sqrt{D_{1}}}\right)^{2}$ (38)
which can also be defined on the density states, generalizes this inner product, because for pure states $\psi_{1}$ and $\psi_{2}$ we have
$F(\psi_{1},\psi_{2})=|\langle\psi_{1}|\psi_{2}\rangle|^{2}$ (39)
but does not give rise to a metric over the entire set of quantum states. But if we consider angles, or equivalently circular arcs, as a measure of the natural distance between pure states in the Bloch model, we are actually using the notion of fidelity to evaluate differences between stimuli.
@@
So, we will use two different notions of distance, one that considers the circular arc between two vector-states, at the surface of the Bloch sphere, and the other one which considers the Euclidean distance between density states, inside the Bloch sphere. The normalized to 1 distance between two such decohered density states
$\begin{pmatrix}\cos^{2}{\theta\over 2}&0\\ 0&\sin^{2}{\theta\over 2}\end{pmatrix}\quad{\rm and}\quad\begin{pmatrix}\cos^{2}{\alpha\over 2}&0\\ 0&\sin^{2}{\alpha\over 2}\end{pmatrix}$ (40)
will be given by (see Figure 4)
$d_{density}(\theta,\alpha)={1\over 2}|\cos\theta-\cos\alpha|$ (41)
Equipped with these two distances, let us now analyse how the phenomenon of categorical perception is naturally expressed in a quantum measurement process. As we mentioned, the points of the sphere’s diameter where the elastic band is stretched, is where the percepts lie, for the measurement in question. These percepts represent an ‘expected reality’, which however is not to be understood as a collection of wild speculations, but rather as the best picture of what is real in relation to the meaning carried by the measured entity, but adapted to the context which the region where the elastic is stretched represents within the mathematical model of the quantum measurement. So the region where the elastic lies does not simply represent an ‘expected reality’, as might be misunderstood from perception theory a la Bruner and Postman, and from the experiment we described, a more correct way to describe what this region represents is the following. ‘That which can be put forward from the context of the measuring apparatus and the cognitive process that takes place as the best hypothesis about reality’. And, as a form of contextual limitation, portions of expected reality creep in to this as a consequence of attempting the best hypothesis about non-contextual reality. Also some contextual biases cannot be avoided, and the warping of categorical perception is one of them. More precisely still, the one who perceives is actively engaged during a measurement process and his or her ‘expected reality’ plays a role in what will occur, together with the reality of the measured entity, expressed by its pure state, which is instead measurement independent. At the level of the contextual region where the elastic lies, the model is Kolmogorovian, the probabilities being an expression of the lack of knowledge of the one who perceives. This is realized in the extended Bloch model by the unpredictable point $\lambda$ where the elastic breaks. Also, the connection between the reality of the considered entity (the pure states describing the stimuli) and the elements of the contextual reality (the on-elastic density states describing the percepts for the given measurement), is illustrated by the deterministic orthogonal fall of the point particle representative of the stimulus onto the elastic band, transforming it into a percept, thus reaching a stage where an outcome (an answer) is actualized, and we are back to a pure state.
@@
To see how a quantum measurement brings about the warping effect of categorical perception, let us consider, to fix ideas, a situation where only two colors exist, Light and Dark, so we are precisely in a situation that can be described in a three-dimensional Bloch sphere. Note that Eleanor Rosch formulated the rationale for the prototype theory for concepts while teaching colors to a primitive community in Papua New Guinea, whose language, called Berinomo, has just two names for colors (Rosch 1973).
@@
Figure 4: We consider a situation where there are two names for colors, which we call Light and Dark, and wish to show that the quantum measurement model, which in this case represents a ‘qubit’, incorporates the phenomenon of categorical perception. For this, we consider three pure states $|\pi/3,\phi\rangle$, $|2\pi/3,\phi\rangle$ and $|0,\phi\rangle$, respectively, in the Bloch representation located in points $A$, $B$ and $C$, which represent stimuli associated with a quantum measurement on this qubit. With each of the three pure states corresponds a density state in which the qubit is located after the measurement, localized in the Bloch representation in respectively points $A^{\prime}$, $B^{\prime}$ and $C^{\prime}$, and described by respectively density matrices (42), (43) and (44). The pure states in $A$ and $B$ belong to two different colors, Light and Dark, and lie at a distance 1/3 from each other. The density states corresponding to them, located in points $A^{\prime}$ and $B^{\prime}$, are at a distance 1/2 from each other. We see here the dilation mechanism of categorical reception at work, for percepts belonging to different categories, Light and Dark. The pure states in $C$ and $A$ belong to the same color, Light, and also lie at a distance 1/3 from each other. The density states corresponding to $C$ and $A$, located in points $C^{\prime}$ and $A^{\prime}$, are at a distance 1/4 from each other. We see here the contraction mechanism of categorical reception at work, for percepts belonging to the same category.

### v2

@@ 3 Quantum Measurement and Categorical Perception
To verify that quantum measurement gives rise to the mechanism of categorical perception, we must be able to express for states how much they differ from each other. But before we get to that, we must identify which states play the role of stimuli and which states play the role of percepts in a quantum measurement process. The stimuli are represented by the states of the quantum entity independent of any measurement that would be going on, and hence they are the pure states. In Hilbert space, these pure states are represented by unit vectors of the Hilbert space. In the Bloch representation, they are represented by the points of the surface of the Bloch sphere. What are the percepts of a quantum measurement?
We must dig deeper into the Hilbert space formalism than what shows us the representation by unit vectors of the pure states to see clearly in it, and the Bloch sphere representation is helpful in this regard. Indeed, the Hilbert space formalism, in addition to the vector space structure that allows us in a simple way to represent the pure states of a quantum entity, and to formulate the superposition principle, there is likewise the density operator formalism, carried by the same Hilbert space, but with its own structure that is further removed from the vector space structure, and is brought out in a less prominent way when quantum mechanics is discussed. Let us first note that the pure states do not coincide with the vectors of Hilbert space but with the unit vectors. In this sense, a superposition is not just a sum, but a sum followed by a renormalization, and is thus in its totality a nonlinear operation although directly using the linear operation, which the sum is. The density operator formalism uses the density operators of the Hilbert space. We will not further specify their mathematical structure at the level of the Hilbert space since they are represented in a more straightforward way by the interior points of the Bloch sphere. Indeed, all interior points of the Bloch sphere represent density states of the considered quantum entity.
It is important to note, and we cover this in depth and in great detail in Aerts et al., (2025), that there is no natural metric for the entire Bloch sphere. There does exist a natural metric on the interior points of the Bloch sphere, that is, on the density states, namely the trace distance. Suppose that $A$ and $B$ are two interior points of the Bloch sphere, the trace distance between $A$ and $B$ is then given by
$T(M(A),M(B))$ $=$ $\frac{1}{2}Tr\left[\sqrt{(M(B)-M(A))^{*}(M(B)-M(A))}\right]$ (1)
@@
where $M(A)$ and $M(B)$ are the density matrices corresponding to the points $A$ and $B$. It can be shown that for a two-dimensional quantum entity, i.e., a qubit, modeled with the Bloch sphere, the trace distance is equal to half the Euclidean distance in the three-dimensional Euclidean space of which the Bloch sphere is a part. It can also be shown that for two pure states $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, hence represented by two points of the surface of the Bloch sphere, the trace distance is given by
@@
This general expression for the trace distance between pure states, makes it clear that the trace distance is not the natural metric to be used between pure states. Indeed, any effect due to quantum coherence disappears when using this metric, since in the general expression (2) the only reference to the states is the inner product between the two states, and hence quantum coherence effects are no longer present.
@@
There is also a purely geometric way by which we can see that the trace distance is unsuitable for measuring distances between pure states. Consider any two pure states, which gives us two points of the surface of the Bloch sphere. If we connect these two points with a line, we see that the points of the line different from these two points always lie in the interior of the Bloch sphere and thus correspond to density states. It is clear that defining a distance associated with a specific geometric entity, the straight line in the three-dimensional Euclidean space to which the Bloch sphere belongs, carries limitations when, as is the case with pure states, this distance is intrinsically intended for elements of a subset with a structure in which the straight line is not contained, and we mean here the spherical surface of the Bloch sphere to which the pure states are restricted.
@@
So we must ask ourselves at this point of our analysis which metric is the proper one to be used on pure quantum states. A natural notion of distance between pure states should only account for the pure states that are in-between the points on the Bloch sphere representing these states, the length of the circular arc connecting these points would be such a quantity. More precisely, if $\theta$ and $\alpha$ are the polar angles of the pure states $|\theta,\phi\rangle$ and $|\alpha,\phi\rangle$ (see Figure 2), the distance to be used to measure how much they are separated from each other, normalized to 1, would then be
$d_{pure}(|\theta,\phi\rangle,|\alpha,\phi\rangle)=\frac{1}{\pi}|\theta-\alpha|$ (3)
@@
Note that it is sufficient to define the distance for two pure states with equal azimuthal angle $\phi$. Indeed, we can always choose a north pole and a south pole for any two pure states such that both states lie in the plane with equal azimuthal angle, so that then the arc of the circle through both points is determined by the difference of the polar angles of both states.
@@
This distance also results from the angle between pure states calculated from their inner product of the Hilbert space, as was identified already in early twentieth century by two mathematicians Guido Fubini and Eduard Study to be the natural metric of the projective space of pure states, and it is now called the Fubini Study metric (Fubini,, 1904; Study,, 1905). For pure states $\psi_{1}$ and $\psi_{2}$ we have
$\gamma(\psi_{1},\psi_{2})=\arccos|\langle\psi_{1}|\psi_{2}\rangle|$ (4)
and normalized it coincides with the arc distance we introduced as natural distance on the pure states in the Bloch sphere.
@@
So, we will use two different notions of distance, one that considers the circular arc between two pure states, at the surface of the Bloch sphere, and the other one which considers the Euclidean distance between density states, inside the Bloch sphere. The normalized to 1 distance between two such density states will be given by (see Figure 2)
$d_{density}(\theta,\alpha)=\frac{1}{2}|\cos\theta-\cos\alpha|$ (5)
Equipped with these two distances, let us now analyze how the phenomenon of categorical perception is naturally expressed in a quantum measurement process.
@@
As we mentioned, the points of the sphere’s diameter between the South Pole and the North Pole are where the percepts lie for the measurement in question. To see how a quantum measurement brings about the warping effect of categorical perception, let us consider, to fix ideas, a situation where only two colors exist, Light and Dark, so we are precisely in a situation that can be described in a three-dimensional Bloch sphere. Note that Eleanor Rosch formulated the rationale for the prototype theory for concepts while teaching colors to a primitive community in Papua New Guinea, whose language, called Berinomo, has only two names for colors (Rosch 1973).
@@
Figure 2: We consider a situation where there are two names for colors, which we call Light and Dark, and wish to show that the quantum measurement model, which in this case represents a ‘qubit’, incorporates the phenomenon of categorical perception. For this, we consider three pure states located in the Bloch representation in points $A$, $B$ and $C$, which represent stimuli associated with a quantum measurement on this qubit. With each of the three pure states corresponds a density state in which the qubit is located after the measurement, in the Bloch representation in respectively points $A^{\prime}$, $B^{\prime}$ and $C^{\prime}$, and described by density matrices. The pure states in $A$ and $B$ belong to two different colors, Light and Dark, and lie at a distance 1/3 from each other. The density states corresponding to them, located in points $A^{\prime}$ and $B^{\prime}$, are at a distance 1/2 from each other. We see here the dilation mechanism of categorical reception at work, for percepts belonging to different categories, Light and Dark. The pure states in $C$ and $A$ belong to the same color, Light, and also lie at a distance 1/3 from each other. The density states corresponding to $C$ and $A$, located in points $C^{\prime}$ and $A^{\prime}$, are at a distance 1/4 from each other. We see here the contraction mechanism of categorical reception at work, for percepts belonging to the same category.